# Non-coherent evolution of closed weakly interacting system leads to equidistribution of probabilities of microstates

### v4

@@ III Non-coherent evolution equation and quantum transitions
We see that if we do not consider the full set of equations (40) describing the evolution of the system’s state through transitions between microstates, but instead focus on the single transition rate (49), we recover the well-known Fermi rule. It is worth noting that the arguments we have presented apply not only to macroscopic closed systems but also, for example, to a single atom emitting a photon into space. The Fermi rule was originally derived for such a situation.
Indeed, we have made two crucial assumptions in our derivation. The first is the applicability of averaging over phases. If the atom and the object with which the emitted photon would interact are not phase-correlated, the second non-coherence mechanism described in the previous section (Fig. 1c) comes into play, allowing averaging to be applicable. The second assumption is the continuous spectrum. While an atom has a discrete spectrum, the system consisting of the atom and the electromagnetic field has a continuous spectrum, so the arguments hold. The key difference between an isolated system and an atom emitting into empty space is that once the photon departs from the atom, no other photons follow, thus disrupting the symmetry of forward and backward transitions.
Since the Bohr model of the atom, the notion of ”quantum jumps” has been present in quantum theory. This is an ironic term for transitions between states of the system that, according to the Bohr model, occurred instantaneously. The theory of decoherence describes ”quantum jumps” as a continuous process of decoherence happening locally on a very short timescale [44]. Our description of quantum transitions allows us to estimate their characteristic timescales.
We have derived the expression (40), where transitions between states are assumed to be instantaneous. However, our derivation allows us to see that the transition time is only negligibly small compared to the characteristic timescale of the evolution of amplitudes. Let us consider the emission of a photon. More generally, we can discuss a transition between fermionic states with the emission of a wave — a boson.
First, we observe that starting from the initial state $|k\rangle$, we can transition equally to two different states: one containing a boson with energy smaller than $\Delta E$, and the other not (Fig. 2b). Thus, such states are ”glued” in our formalism. Hence, bosons with energies less than $\Delta E$ are practically indistinguishable from the absence of any bosons. On a timescale $\Delta t$, we can only ”detect” bosons with energy $\hbar\omega\geq\hbar/\Delta t$, meaning $T\leq\Delta t$, where $T$ is the wave period. On the other hand, we cannot separate events in time that are separated by less than $\Delta t$. The transition time is the minimal time interval that still allows the detection of the emission: $\tau=\mathrm{min}{\Delta t:T\leq\Delta t}$. This implies $\tau=T$; the transition time is equal to the period of the emitted wave, which makes perfect sense.
In the next section, we will suggest a way to include the waves of small frequency within the presented framework.
@@ IV Applications to statistical mechanics
We want to use the derived equation to address the problem of irreversibility in statistical mechanics. We consider all three instances of this problem mentioned in the introduction: the second law of thermodynamics, the equidistribution of states in equilibrium, and the irreversibility of the Boltzmann equation. In [14], it is proven that entropy, defined as
@@
increases in a system whose evolution is described by the Fermi master equation, unless it has already reached its maximum. Entropy, defined by equation (50), attains its maximum when the probabilities of all states are equal. Thus, this also proves the equiprobability of states in equilibrium. However, we can approach the problem from the opposite direction.
@@
Now we want to derive the Boltzmann collision integral. In the second section, we have described two mechanisms leading to the lack of coherence in interaction. Here we consider two particular types of interactions: the emission and absorption of phonons by electrons and the three phonons interaction. In the third section we found that for our theory to be applicable, the mean time of quantum transitions should be much higher than the coherence time: $t_{\text{coh}}/T_{A}\ll 1$. If we switch from times to lengths, the mean free passes of particles are much higher than the coherence lengths. If also the size of the system is much higher than the mean free paths of the particles, most interactions in the systems occur between particles whose emissions are not causally related. In the first example, a phonon is typically absorbed by a different electron than the one that emitted it. In the second example, the interaction of phonons occurs between phonons that were emitted at a distance much greater than the coherence length from each other. It gives us right to perform the averaging over phases thus leading to the applicability of the non-coherent evolution formula (40).
Let us consider the first case: the transitions between energy levels by electrons with the emission or absorption of phonons. More generally, it could be any fermions and bosons. The interaction has the form $V_{ij}\hat{c}^{\dagger}_{i}\hat{c}_{j}\hat{a}$, where $\hat{c},\hat{a}$ are annihilation operators for fermions and bosons, respectively, and indices $i,j$ denote fermionic single-particle states.
@@
Thus, we have derived two cases of the Boltzmann collision integrals. With this derivation, the irreversibility of the Boltzmann equations comes as no surprise, since the irreversibility of the collision part is inherited from the irreversibility of the non-coherent evolution equation (40), from which it is derived.
During the derivation of the non-coherent evolution equation (40), we noted that we should pick $\Delta t$ small compared to the typical timescale of the change of probability amplitude: $\Delta t\ll T_{A}$ (Fig. 2). At the end of the previous section, we observed that the period of emitted waves in a given transition, $T_{ij}$, should be less than $\Delta t$, which yields $T_{ij}\ll T_{A}=\hbar/V_{ij}$. We can see that for long waves with period $T_{ij}\to\infty$, this condition cannot be fulfilled unless the matrix element of interaction with long-wavelength bosons, $V_{ij}$, tends to zero as the wavelength goes to zero. Additionally, in the previous paragraph, we noted that low-frequency waves were excluded from our derivations. We suggest that low-frequency waves contribute to the force term in the Boltzmann equation and, therefore, should be treated separately. In this paper, we have derived the interaction in a local region, which corresponds to the collision term in the Boltzmann equation. The free movement of particles in the media, corresponding to the diffusion term, and interaction with low-frequency fields, corresponding to the force term, should be derived separately. One possibility is the approach described in Ref. [7].
@@
V Summary and conclusions
The present work develops a theoretical framework connecting quantum mechanics with irreversible statistical processes through the concept of non-coherent evolution. We have demonstrated how averaging over phases in quantum systems leads to the emergence of irreversible behavior, providing a bridge between reversible quantum mechanics and irreversible statistical mechanics. We also provided an explanation of how this averaging arises.
We can loosely divide the manuscript into two parts. In the first part (Section 2 and most of Section 3), we derive the equation of non-coherent evolution. In the second part (the end of Section 3 and Section 4), we use it to resolve problems associated with our understanding of quantum transitions and the irreversible behaviour of macroscopic systems.
At the beginning, we considered non-coherent transformations, showing how they naturally arise from unitary transformations when phase information is lost. Non-coherence arises from the fact that no wave has exactly one frequency but is composed of a sheaf of waves with closely spaced frequencies. Such sheaves can be mutually coherent or non-coherent. We identified two fundamental mechanisms leading to the loss of coherence between sheaves: temporal separation exceeding the coherence time and interaction between causally unrelated emissions. These mechanisms were illustrated using optical systems, where direct experimental verification is possible.
Then we move on from studying a discrete series of transformations to continuous evolution. We derive the non-coherent evolution equation, which is the central result of this work ((45), (40)). Our analysis begins with partitioning the wave function of the system into two components: fast-varying psi functions of microstates and probability amplitudes, which are smooth (in-time) envelopes of the psi function (21). Our derivation relies on the assumption of weak interaction and the existence of well-separated time scales. We introduced a technical timescale $\Delta t$, chosen between the microstate oscillation period and the characteristic scale of amplitude variation. Using this technical assumption, we showed that for timescales much larger than the coherence time, unitary evolution transforms into symmetric stochastic evolution, described by continuous-time Markov chains. We then show that the final results of the calculations do not depend on the value of the timescale $\Delta t$.
@@ V Summary and conclusions
There is also one dubious point in the classical derivation of Fermi’s rule: we integrate over an infinite time period to obtain a momentary transition rate. The introduction of the technical timescale makes this point very clear. This timescale is very large – ”infinite” – compared to the microstate oscillation period, and very small – ”infinitesimal” – compared to the characteristic scale of amplitude variation. Thus, we can treat it accordingly at different points in the derivation. Our approach also permits us to estimate the time interval of a quantum transition. It turns out that this time coincides with the period of the wave that is being emitted. Thus, we believe that the presented theory deepens our understanding of quantum transitions.
@@
Our results suggest that the physical concept of non-coherence and the mathematical structure of symmetric continuous-time Markov chains are fundamental to understanding the emergence of irreversibility in quantum systems, potentially extending beyond the weak-interaction regime considered in the present manuscript.
@@ Acknowledgements.
The author would like to thank Miguel Larotonda, who patiently answered my questions about coherence in the context of classical and quantum optics, and whose insightful responses greatly improved my understanding of these areas.

### v5

@@ III Non-coherent evolution
We see that if we do not consider the full set of equations (40) describing the evolution of the system’s state through transitions between microstates, but instead focus on the single transition rate (49), we recover the well-known Fermi rule. Thus, we re-derived the formula for calculating transition rates between states.
Since the Bohr model of the atom, the notion of quantum jumps has been part of quantum theory. The term is somewhat ironic, as in the Bohr model these transitions between states were assumed to occur instantaneously. In decoherence theory, by contrast, quantum jumps are understood as a continuous process of local decoherence occurring on a very short timescale [44]. In our derivation, transitions appear instantaneous only when compared with the technical timescale $\Delta t$. In principle, the techniques developed here could allow one to estimate the actual transition time, though doing so would require a different set of ideas than those presented in this manuscript.
@@ IV Applications to statistical mechanics
We want to use the derived equation to address the problem of irreversibility in statistical mechanics. We consider all three instances of this problem mentioned in the introduction: the second law of thermodynamics, the equidistribution of states in equilibrium, and the irreversibility of the Boltzmann equation.
For the secon law we use the Gibbs entropy defined as
@@
We note that we do not use von Neumann entropy or other types of quantum entropies [16]. The von Neumann entropy has the advantage of being invariant under transformations of basis. However, in our treatment, as we explained at the end of Section II we already work in a specified basis. Also, the goal of our manuscript is to derive the Gibbs statistical mechanics, so we use Gibbs entropy.
In [14], it is proven that entropy defined by formula (50) increases in a system whose evolution is described by the Fermi master equation, unless it has already reached its maximum. Gibbs entropy attains its maximum when the probabilities of all states are equal. Thus, this also proves the equiprobability of states in equilibrium. However, we can approach the problem from the opposite direction.
@@
Now we want to derive the Boltzmann collision integral for two particular cases. Let us consider the first case: the transitions between energy levels by electrons with the emission or absorption of phonons. More generally, it could be any fermions and bosons. The interaction has the form $V_{ij}\hat{c}^{\dagger}_{i}\hat{c}_{j}\hat{a}$, where $\hat{c},\hat{a}$ are annihilation operators for fermions and bosons, respectively, and indices $i,j$ denote fermionic single-particle states.
@@
Thus, we have derived two cases of the Boltzmann collision integral. In this derivation, the irreversibility of the Boltzmann equation is no surprise, since the irreversibility of the collision term is inherited from the irreversibility of the non-coherent evolution equation (40), from which it is derived. This stands in sharp contrast to the BBGKY approach to the derivation of the Boltzmann equation [6]. In BBGKY, irreversibility emerges as a consequence of the transition from many-particle to single-particle distribution functions, whereas in the present theory the evolution of the many-particle distribution function is already irreversible.
After performing the smoothing operation (25), only contributions from waves with frequencies much larger than $1/\Delta t$ are retained, while information about low-frequency components is entirely suppressed. We propose that such low-frequency modes contribute to the force term in the Boltzmann equation and should therefore be treated separately. In the present work, we have derived the interaction within a local region, which corresponds to the collision term in the Boltzmann equation. By contrast, the free propagation of particles, corresponding to the diffusion term, and the coupling to low-frequency fields, corresponding to the force term, must be obtained through an independent treatment.
To this end, we suggest extending the formalism by allowing the amplitudes of the psi-function (21) to depend slowly on the spatial coordinates, in addition to their slow temporal variation. Under this assumption, the non-collision contributions to the Boltzmann equation can be derived consistently, with low-frequency fields entering explicitly as a force term. In the present paper, however, we restrict attention to the homogeneous case, where the amplitudes are taken to be spatially uniform and external low-frequency fields are absent. This is the precise meaning of ’homogeneity’ as introduced in the list of assumptions at the beginning of Section III.
@@
V Summary and discussion
The present work develops a theoretical framework connecting quantum mechanics with irreversible statistical processes through the concept of non-coherent evolution. We have demonstrated how averaging over phases in quantum systems leads to the emergence of irreversible behavior, providing a bridge between reversible quantum mechanics and irreversible physical kinetics. We also provided an explanation of how this averaging arises. The most important feature of this paper is the explanation of irreversible behaviour in the isolated system, which explains the second law of thermodynamics and the microcanonical ensemble as an equilibrium state of the system.
We can loosely divide the manuscript into two parts. In the first part (Section 2 and Section 3), we derive the equation of non-coherent evolution. In the second part (Section 4), we use it to resolve problems associated with the irreversible behaviour of macroscopic systems.
At the beginning, we considered non-coherent transformations, showing how they naturally arise from unitary transformations when phase information is lost. Non-coherence arises from the fact that no wave has exactly one frequency but is composed of a sheaf of waves with closely spaced frequencies. Such sheaves can be mutually coherent or non-coherent. These mechanism was illustrated using optical systems, where direct experimental verification is possible.
Then we move on from studying a discrete series of transformations to continuous evolution. We derive the non-coherent evolution equation, which is the central result of this work (40, 45). Our analysis begins with partitioning the wave function of the system into two components: fast-varying psi functions of microstates and probability amplitudes, which are smooth (in-time) envelopes of the psi function (21). Our derivation relies on the assumption of weak interaction and the existence of well-separated time scales. We introduced a technical timescale $\Delta t$, chosen between the microstate oscillation period and the characteristic scale of amplitude variation. Using this technical assumption, we showed that for timescales much larger than the coherence time, unitary evolution transforms into symmetric stochastic evolution, described by continuous-time Markov chains. We then show that the final results of the calculations do not depend on the value of the timescale $\Delta t$.
@@ V Summary and discussion
There is also one dubious point in the classical derivation of Fermi’s rule: we integrate over an infinite time period to obtain a momentary transition rate. The introduction of the technical timescale makes this point very clear. This timescale is very large – ”infinite” – compared to the microstate oscillation period, and very small – ”infinitesimal” – compared to the characteristic scale of amplitude variation. We can treat it accordingly at different points in the derivation. Thus, we believe that the presented theory deepens our understanding of quantum transitions.
@@
In the introduction, we briefly outlined existing theories of irreversibility and indicated how the approach in this manuscript differs from them. Now that we have presented our findings, let us revisit this comparison in more detail.
First, consider the measurement problem and its treatment in decoherence theory. While our theory does not address the problem of outcomes, we argue that no definite outcomes are required to derive statistical mechanics from quantum mechanics. This is because only expectation values are needed, and in statistical mechanics we deal with very large systems, where the law of large numbers guarantees that actual outcomes do not differ significantly from the expected values.
Furthermore, since the early days of quantum theory, it has been suggested that measurement results from the interaction of a quantum system with a macroscopic body. We understand a macroscopic body as one to which the laws of classical statistical physics, as developed by Gibbs and Boltzmann, can be applied. We hope that clarifying what constitutes a macroscopic body, without appealing to interactions with the environment, will allow us to escape the vicious circle: the definition of macroscopic bodies through decoherence, and the definition of decoherence as the result of interaction with a macroscopic body. A clearer understanding of what qualifies as a macroscopic body may then shed light on what happens when a coherent quantum system interacts with it, and thus open a new perspective on the measurement problem.
Second, let us turn to quantum thermodynamics. While it is true that the equation of non-coherent evolution (40, 45) is in many respects analogous to those found in quantum thermodynamics, we hope it is clear that both the physical picture offered by the present theory and the mathematical methods employed are very different. Concepts such as coherence length are not introduced in quantum thermodynamics. We may hope that further development of the theory will provide a testable and decisive distinction between the two approaches.
We have already emphasized the distinction of the present approach, which treats non-coherence as an intrinsic property of the system rather than as a consequence of interaction with the environment. Shifting the focus away from interaction is the most radical departure from almost all existing theories – not only those based on quantum mechanics and decoherence, but even those rooted in classical mechanics, such as BBGKY hierarchy [6]. All the mentioned theories in their core attribute irreversibility to interaction. Indeed, the theories based on classical mechanics, roughly speaking, attribute irreversibility to chaotization, which is a result of interatomic collisions. In our theory, in contrast, interaction is not a cause of irreversibility.
If it were not for non-coherence, interactions such as the emission of a boson by a fermion (used as an example throughout this text) would occur reversibly, as described by equation (30). A single interaction would place the system into a superposition of states with and without the emitted boson. It is in direct analogy with our optical experiment in coherent regime, where photons after an incidence at an interface are placed into a superposition of being reflected and transmitted (Fig. 2b). The full set of all interactions would then yield a superposition of many microstates, each with a specific amplitude. Such a giant superposition evolves unitarily, according to (30), and is therefore reversible, even when all interactions are included.
It is only when we impose the condition of non-coherence that the situation changes. The system then reduces to one in which all microstates have definite probabilities but lack definite phases, analogous to photons in the non-coherent experiment (Fig. 2c). The evolution of such a system is governed by the stochastic equation (40, 45), which is irreversible. Thus, in the present viewpoint, it is not interaction but non-coherence that constitutes the true origin of irreversibility.
Now, how can we test the theory? After all, in this paper we started from a well-known equation (the Schrödinger equation) and, after some derivations, arrived at well-known results (the microcanonical ensemble, the Boltzmann collision integral). The experiment presented in Ref. [37] does serve as a test: it shows that light, traditionally described by reversible wave equations, exhibits irreversible properties due to non-coherence. Yet this experiment is probably too simple to be fully convincing.
One possible way of testing the theory would be, in some sense, the opposite experiment: namely, to identify wave-like behavior in transport phenomena and, crucially, to devise a method for measuring the coherence length in such systems. We suggest that transport in nanostructures offers a promising opportunity, since the characteristic scale of these systems may be smaller than or comparable to the coherence length. The essential task would be to vary system parameters in order to observe the transition from coherent to non-coherent transport, under conditions where the particle mean free path remains much larger than the structural scale. Demonstrating such a transition would establish that the mean free path and the coherence length are distinct physical quantities, thereby providing unambiguous evidence that the loss of coherence cannot be attributed to interactions.
Transport in nanostructures is a rapidly developing field [46], and the distinction between coherent and non-coherent transmission has already been discussed in relation to interfaces [47, 48]. It is therefore quite possible that some effects consistent with the present theory are already reported in the literature. Otherwise, we may hope to design a realistic experiment to verify the theory in the near future.
From a theoretical perspective, an important investigation direction is to study systems that are neither fully coherent nor fully non-coherent, systems in which the coherence length must explicitly enter the results of calculations. This is equally relevant in the context of optics and in the context of transport theories at the nanoscale. While coherence forms a spectrum, ranging from fully coherent through partially coherent to non-coherent (see Eq. 11), reversibility is a binary property: a process is either reversible or irreversible. To fully understand irreversibility, one must analyze the (ir)reversible behavior of systems in which coherence is only partially lost.
Another important direction is to extend the scope of the theory, as was briefly outlined in the manuscript: namely, to incorporate perturbations of energies and transition rates due to interactions, and to generalize the framework to non-homogeneous cases.
@@ Acknowledgements.
The author would like to thank Miguel Larotonda, who patiently answered my questions about coherence in the context of classical and quantum optics, and whose insightful responses greatly improved my understanding of these areas.During the preparation of this work, the author used ChatGPT in order to improve the quality of English. After using this service, the author reviewed and edited the content as needed and takes full responsibility for the content of the publication.